A project will produce an operating cash flow of $315,000 a year for four years. The...

  1. A project will produce an operating cash flow of $315,000 a year for four years. The initial cash outlay for equipment will be $874,000. The net aftertax salvage value of $43,000 will be received at the end of the project. The project requires 78,000 of net working capital up front that will be fully recovered. What is the net present value of the project if the required rate of return is 13 percent?

    $74,311.68

    $69,893.52

    $64,410.27

    $59,170.03

    $54,736.08
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=315,000/1.13+315,000/1.13^2+315,000/1.13^3+315,000/1.13^4+43000/1.13^4+78000/1.13^4

=1011170.03

Present value of outflows=874,000+78000

=952000

NPV=Present value of inflows-Present value of outflows

=1011170.03-952000

=$59170.03(Approx).